Katherine saved $8 on a CD player due to a 20% discount. What was the original price of the CD player?
Answer by stanbon(75887)

Katherine saved $8 on a CD player due to a 20% discount. What was the original price of the CD player?
------
Let the original price be "x".
----
0.20x = 8
x = 8/0.2
x = 8/(1/5) = $40
---------------------
Cheers,
Stan H.
-------------